# Joint Transmit and Pinching Beamforming for Pinching Antenna Systems (PASS): Optimization-Based or Learning-Based?

### v2

@@ Abstract
A novel pinching antenna system (PASS)-enabled downlink multi-user multiple-input single-output (MISO) framework is proposed. PASS consists of multiple waveguides spanning over thousands of wavelength, which equip numerous low-cost dielectric particles, named pinching antennas (PAs), to radiate signals into free space. The positions of PAs can be reconfigured to change both the large-scale path losses and phases of signals, thus facilitating the novel pinching beamforming design. A sum rate maximization problem is formulated, which jointly optimizes the transmit and pinching beamforming to adaptively achieve constructive signal enhancement and destructive interference mitigation. To solve this highly coupled and nonconvex problem, both optimization-based and learning-based methods are proposed. 1) For the optimization-based method, a majorization-minimization and penalty dual decomposition (MM-PDD) algorithm is developed, which handles the nonconvex complex exponential component using a Lipschitz surrogate function and then invokes PDD for problem decoupling. 2) For the learning-based method, a novel Karush-Kuhn-Tucker (KKT)-guided dual learning (KDL) approach is proposed, which enables KKT solutions to be reconstructed in a data-driven manner by learning dual variables. Following this idea, a KDL-Tranformer algorithm is developed, which captures both inter-PA/inter-user dependencies and channel-state-information (CSI)-beamforming dependencies by attention mechanisms. Simulation results demonstrate that: i) The proposed PASS framework significantly outperforms conventional massive multiple input multiple output (MIMO) system even with a few PAs. ii) The proposed KDL-Transformer can improve over $30\%$ system performance than MM-PDD algorithm, while achieving a millisecond-level response on modern GPUs.
Index Terms:
Beamforming, deep learning, flexible antennas, pinching antenna system (PASS), pinching beamforming.
@@ I Introduction
Shannon’s theorem establishes the capacity upper bound of wireless communication systems based on the signal-to-noise (SNR) ratio [1], which dramatically depends on channel qualities of mobile users. To break through the limitation of fixed-channel assumptions and make radio environment customizable, recent advances in sixth-generation (6G) wireless networks have been devoted to developing several new forms of flexible-antenna techniques, such as reconfigurable intelligent surfaces (RISs) [2, 3], fluid antennas [4], and movable antennas [5]. Specifically, RISs equip a large amount of low-cost passive reflective/transmissive meta-material elements to reconfigure the communication environment by altering the phase shifts and amplitudes of incident signals [2, 3]. More recently, fluid antenna and movable antenna systems have been proposed [4, 5], which enable the transceivers/receivers to dynamically adjust antenna positions and their electromagnetic properties, thus creating favorable spatial channels for multi-input multi-output (MIMO) communications.
Despite the promises, there are two limitations of existing flexible-antenna techniques. On the one hand, the transceivers/receivers typically spans across a restricted spatial range, which significantly limits their capabilities to reconfigure large-scale path loss and maintain stable line-of-sight (LoS) links for mobile users. Particularly, in RISs-aided systems, the received signals typically undergo the channel fading of both the base station (BS)-RIS links and the RIS-user links. Moreover, fluid antennas and movable antennas usually move within a few wavelengths, which lead to minor effects on the large-scale path loss and suffer from susceptibility to barrier blockage in high-frequency communications. On the other hand, existing flexible-antenna techniques are usually equipped with a fixed number of antennas, which are difficult to flexibly scale up/down according to realistic demands.
@@
To circumvent these drawbacks, pinching antenna system (PASS) has recently emerged as a novel flexible-antenna technique [6, 7]. PASS is built on a rod-like transmission line, named dielectric waveguide, whose length spans from a few meters to tens of meters, and even up to thousands of times longer than the wavelength [8]. The dielectric waveguide acts as a leaky-wave antenna, where multiple radiation points, implemented by dielectric particles, can be flexibly pinched and released anywhere along this waveguide. These small dielectric particles are known as pinching antennas (PAs), which enables radio-wave transmission and reception of PASS. Amongst existing flexible antenna techniques, PASS has several unique advantages [7]: 1) Path-loss reconfiguration and LoS link construction: PASS can reconfigure the large-scale path loss to deliver wireless services to the last-meter range, and maintain stable LoS links to avoid blockage in high-frequency communications. 2) Flexible beamforming: The deployment of PAs can transform the entire electromagnetic propagation environment, thus achieving flexible beam focusing/steering effects even with a few PAs. 3) Scalability: PAs can be easily activated, shifted, and released from the dielectric waveguides to accommodate various traffic demands.
@@
  • The deployment of each PA impacts both the large-scale path loss and the phases of signals, which leads to the novel concept of pinching beamforming. Despite providing additional degree-of-freedom (DoF), efficient pinching beamforming is required to simultaneously reconfigure large-scale fading and signal phases, which is more challenging than conventional MIMO beamforming and existing flexible-antenna techniques.
@@
To address above issues, this paper proposes a novel PASS-enabled downlink multi-user MISO framework. A sum rate maximization problem is formulated by jointly optimizing the newly introduced pinching beamforming and the conventional transmit beamforming. The resulting optimization problem is highly coupled and involves nonconvex complex exponential components. To tackle this problem, this paper proposes both optimization-based and learning-based methods. First, by reformulating the coupled optimization problem using the weighted minimum mean square error (WMMSE), we develop a majorization-minimization and penalty dual decomposition (MM-PDD) algorithm, which is guaranteed to find the stationary solution. Specifically, the Lipschitz gradient surrogate function is constructed for convex relaxation of complex exponential components. Then, the coupled optimization is decomposed by PDD theory and iteratively solved by block coordinate descent (BCD). Secondly, to achieve fast beamforming with a millisecond-level response, we propose a novel optimization theory-guided learning method, termed Karush-Kuhn-Tucker (KKT)-guided dual learning (KDL). The key idea is to enable machine learning (ML) to directly reconstruct Karush-Kuhn-Tucker (KKT) solutions by learning only a few dual variables in a data-driven manner. Hence, it inherits the white-box KKT solution structure, but avoids inefficient alternating block descents required by iterative optimization algorithms. We implement KDL by exploiting the powerful large language model (LLM) techniques, and comes up with a KDL-Transformer algorithm. The channel-state-information-(CSI)-to-beamforming mapping is regarded as a sequence-to-sequence learning task. By modelling both inter-PA/inter-user dependencies and CSI-beamforming dependencies, the proposed KDL-Transformer achieves stunning performance gains compared to mathematical optimization and black-box learning algorithms.
The main contributions of this work are summarized as follows.
  • We propose a novel PASS-enabled downlink multi-user MISO framework, which enables the additional pinching beamforming by configuring PAs’ locations. A joint transmit and pinching beamforming optimization problem is formulated to maximize the system sum rate achieved by PASS. To tackle this coupled nonconvex problem, we propose both optimization-based and learning-based methods.
  • For the optimization-based method, we develop an MM-PDD algorithm. The nonconvex coupled optimization problem is first reformulated into a WMMSE problem. We derive Lipschitz gradient surrogate function based on MM to handle the nonconvex complex exponential component, and then decouple the problem based on PDD to obtain stationary solutions.
@@
  • We provide numerical results to demonstrate the effectiveness of the proposed framework and algorithms, which demonstrates that: 1) The proposed PASS architecture outperforms conventional hybrid beamforming architecture even only a few PAs are activated. 2) Unlike black-box learning algorithms that are stuck at inferior points, the proposed KDL-Transformer increases over $30\%$ system performance than MM-PDD algorithm, while achieving over $7000$ times faster execution on modern GPUs.
@@
Notations: The variable, vector, and matrix are denoted by $x$, $\mathbf{x}$, and $\mathbf{X}$, respectively. $|x|$ denotes the absolute value of a real number and the modulus of a complex number. $\mathrm{Re}\left\{x\right\}$ and $\mathrm{Im}\left\{x\right\}$ denote the real and image parts of $x$, and $x^{H}$ is the complex conjugate number of $x$. $\mathbf{1}_{N\times 1}$ denotes an $N$-dimension all-ones vector. $\mathbf{X}^{T}$ and $\mathbf{X}^{H}$ denote the transpose and the Hermitian matrix. $\left\|\mathbf{x}\right\|_{\infty}=\max_{i}\left|x_{i}\right|$, $\left\|\mathbf{x}\right\|$ is the vector Euclidean norm, and $\left\|\mathbf{X}\right\|$ is the matrix Frobenius norm.
@@ II System Model and Problem Formulation
In this paper, we propose a PASS-enabled downlink multi-user MISO framework, as shown in Fig. 1. The BS is equipped with $N$ waveguides and $M$ PAs to serve $K$ single-antenna users. Each waveguide stretches thousands of wavelengths. There are $L$ PAs flexibly pinched along each waveguide, and the total number of PAs is $M=L\times N$. By activating PAs at different points along the waveguides, both the phases of incident signals and the large-scale fading can be altered. Let $\mathcal{N}$ and $\mathcal{M}=\mathcal{L}_{1}\cup\mathcal{L}_{2}\cup\dots\mathcal{L}_{N}$ denote the sets of all the waveguides and PAs, respectively, where $\mathcal{L}_{n}=\left\{a_{n,1},a_{n,2},\dots,a_{n,L}\right\}$ is the collection of PAs at waveguide $n$. To support spatial multiplexing, each waveguide is connected to a dedicated radio frequency (RF) chain, which converts the signal multiplexed at the baseband and feeds it into the waveguide. To ensure multiplexing gains, we assume $N=K$. Both waveguides and PAs are installed at a fixed height of $h^{\mathrm{PA}}$, and the resulting PASS spans across a rectangular area with a size of $S_{\mathrm{x}}\times S_{\mathrm{y}}$ $\text{m}^{2}$. The three-dimension Cartesian coordinate of the feed point for the $n$-th waveguide is given by $\bm{\eta}_{n}^{\mathrm{W}}=\left[0,y_{n}^{\mathrm{W}},h^{\mathrm{PA}}\right]^{T}$, where $y_{n}^{\mathrm{W}}$ is the $y$-axis coordinate of this waveguide. The location of each PA $a_{n,l}$ along the $n$-th waveguide can be defined as $\bm{\eta}_{n,l}\left(x_{n,l}\right)=\left[x_{n,l},y_{n,l},h^{\mathrm{PA}}\right]^{T}$, where $x_{n,l}$ is the adjustable pinched location of PA $a_{n,l}$ along $x$-axis, and $y_{n,l}$ is the fixed coordinate over $y$-axis that depends on the connected waveguide, i.e., $y_{n,l}=y_{n}^{\mathrm{W}}$. Let $\mathbf{x}_{n}=\left[x_{n,1},x_{n,2},\dots,x_{n,L}\right]^{T}\in\mathbb{R}^{L\times 1}$ stack the $x$-axis locations of PAs over the $n$-the waveguide, which satisfy $0\leqslant x_{n,1}<x_{n,2}<\dots<x_{n,L}\leqslant S_{\mathrm{x}}$, $\forall n\in\mathcal{N}$. $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\right]\in\mathbb{R}^{N\times L}$ stacks the locations of all PAs. Furthermore, the location of user $k$ is denoted by $\bm{\eta}_{k}^{\mathrm{U}}=\left[x_{k}^{\mathrm{U}},y_{k}^{\mathrm{U}},0\right]^{T}$, $\forall k\in\mathcal{K}$.
@@ II-A Signal Radiation Model
where $\lambda_{\mathrm{W}}=\frac{\lambda_{f}}{n_{\mathrm{eff}}}$ is the guided wavelength, where $n_{\mathrm{eff}}$ denoting the effective refractive index of the dielectric waveguide [14, 6]. Moreover, $\left\|\bm{\eta}_{n,l}\left(x_{n,l}\right)-\bm{\eta}_{n}^{\mathrm{W}}\right\|=x_{n,l}$ denotes the distance from waveguide $n$ to PA $a_{n,l}$. Since the transmitting power is equally radiated by each PA along the waveguide, the amplitude of the transmitted signal is multiplied by $1/\sqrt{L}$.
@@
Let $\mathbf{h}_{k}^{H}\left(\mathbf{X}\right)\!=\!\left[\mathbf{h}_{k,1}^{H}\!\left(\mathbf{x}_{1}\right),\mathbf{h}_{k,2}^{H}\!\left(\mathbf{x}_{2}\right),\dots,\mathbf{h}_{k,N}^{H}\!\left(\mathbf{x}_{N}\right)\right]\in\mathbb{C}^{1\times M}$ stack the channel vectors from all the PAs to user $k$. Hence, the received signal at user $k$ through the channels manipulated by PAs can be compactly expressed by
[…]
@@
$\gamma_{k}\!=\!\frac{\left|\mathbf{h}_{k}^{H}\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k}\right|^{2}}{\sum\limits_{k^{\prime}\neq k}\!\left|\mathbf{h}_{k}^{H}\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k^{\prime}}\right|^{2}\!\!+\!\sigma^{2}}\!\!=\!\!\frac{E_{k}}{\sum\limits_{k^{\prime}\neq k}I_{k,k^{\prime}}\!+\!\sigma^{2}}.$ (9)

### v3

# Joint Transmit and Pinching Beamforming for Pinching Antenna System (PASS): Optimization-Based or Learning-Based?

@@ Abstract
A novel pinching antenna system (PASS)-enabled downlink multi-user multiple-input single-output (MISO) framework is proposed. PASS consists of multiple waveguides spanning over thousands of wavelength, which equip numerous low-cost dielectric particles, named pinching antennas (PAs), to radiate signals into free space. The positions of PAs can be reconfigured to change both the large-scale path losses and phases of signals, thus facilitating the novel pinching beamforming design. A sum rate maximization problem is formulated, which jointly optimizes the transmit and pinching beamforming to adaptively achieve constructive signal enhancement and destructive interference mitigation. To solve this highly coupled and nonconvex problem, both optimization-based and learning-based methods are proposed. 1) For the optimization-based method, a majorization-minimization and penalty dual decomposition (MM-PDD) algorithm is developed, which handles the nonconvex complex exponential component using a Lipschitz surrogate function and then invokes PDD for problem decoupling. 2) For the learning-based method, a novel Karush-Kuhn-Tucker (KKT)-guided dual learning (KDL) approach is proposed, which enables KKT solutions to be reconstructed in a data-driven manner by learning dual variables. Following this idea, a KDL-Transformer algorithm is developed, which captures both inter-PA/inter-user dependencies and channel-state-information (CSI)-beamforming dependencies by attention mechanisms. Simulation results demonstrate that: i) The proposed PASS framework significantly outperforms conventional massive multiple input multiple output (MIMO) system even with a few PAs. ii) The proposed KDL-Transformer can improve over $20\%$ system performance than MM-PDD algorithm, while achieving a millisecond-level response on modern GPUs.
@@ I Introduction
Shannon’s theorem establishes the capacity upper bound of wireless communication systems based on the signal-to-noise ratio (SNR) [1], which dramatically depends on channel qualities of mobile users. To break through the limitation of fixed-channel assumptions and make radio environment customizable, recent advances in sixth-generation (6G) wireless networks have been devoted to developing several new forms of flexible-antenna techniques, such as reconfigurable intelligent surfaces (RISs) [2, 3], fluid antennas [4], and movable antennas [5]. Specifically, RISs equip a large amount of low-cost passive reflective/transmissive meta-material elements to reconfigure the communication environment by altering the phase shifts and amplitudes of incident signals [2, 3]. More recently, fluid antenna and movable antenna systems have been proposed [4, 5], which enable the transceivers to dynamically adjust antenna positions and their electromagnetic properties, thus creating favorable spatial channels for multi-input multi-output (MIMO) communications.
@@
Existing flexible-antenna technologies are effective in shaping the scattering environment to improve channel conditions. However, since conventional MIMO transceivers typically operate within a limited spatial range, their ability to influence large-scale fading remains restricted. Recently, the pinching antenna system (PASS) has emerged as a novel flexible-antenna technique [6, 7], which enables direct manipulation of path loss. PASS is built on a rod-like transmission line, known as a dielectric waveguide, that can span from a few meters to tens of meters, or even up to thousands of times longer than the wavelength [8]. The dielectric waveguide acts as a leaky-wave antenna, along which multiple radiation points, termed pinching antennas (PAs), can be flexibly pinched and released at arbitrary locations. These PAs are implemented by small dielectric particles non-contact to the waveguides. Different from conventional leaky-wave systems with fixed radiating elements, PASS enables dynamic and reconfigurable activation and placement of these PAs in response to varying user and channel conditions. This unique design of PASS leads to several advantages [7]: 1) Path-loss reconfiguration and LoS link construction: PASS can reconfigure the large-scale path loss to deliver wireless services to the last-meter range, and maintain stable line-of-sight (LoS) links to avoid blockage in high-frequency communications. 2) Flexible beamforming: The deployment of PAs can transform the entire electromagnetic propagation environment, thus achieving flexible beam focusing/steering effects even with a few PAs. 3) Scalability: PAs can be easily activated, shifted, and released from the dielectric waveguides to accommodate various traffic demands.
@@
  • The deployment of each PA impacts both the large-scale path loss and the phases of signals, which leads to the novel concept of pinching beamforming. Despite providing additional degree-of-freedom (DoF), efficient pinching beamforming is required to simultaneously alter large-scale fading and signal phases for both path loss reduction and multi-user interference mitigation, which is challenging even in a LoS-dominant environment.
@@
To address above issues, this paper proposes a novel PASS-enabled downlink multi-user MISO framework. A sum rate maximization problem is formulated by jointly optimizing the newly introduced pinching beamforming and the conventional transmit beamforming. The resulting optimization problem is highly coupled and involves nonconvex complex exponential components. To tackle this problem, this paper proposes both optimization-based and learning-based methods. First, by reformulating the coupled optimization problem using the weighted minimum mean square error (WMMSE), we develop a majorization-minimization and penalty dual decomposition (MM-PDD) algorithm, which is guaranteed to find the stationary solution. Specifically, the Lipschitz gradient surrogate function is constructed for convex relaxation of complex exponential components. Then, the coupled optimization is decomposed by PDD theory and iteratively solved by block coordinate descent (BCD). Secondly, to achieve fast beamforming with a millisecond-level response, we propose a novel optimization theory-guided learning method, termed Karush-Kuhn-Tucker (KKT)-guided dual learning (KDL). The key idea is to enable machine learning (ML) to directly reconstruct KKT solutions by learning only a few dual variables in a data-driven manner. Hence, it inherits the white-box KKT solution structure, but avoids inefficient alternating block descents required by iterative optimization algorithms. We implement KDL by exploiting the powerful large language model (LLM) techniques, and comes up with a KDL-Transformer algorithm. By casting the channel-state-information-(CSI)-to-beamforming mapping as a sequence-to-sequence learning task, both inter-PA/inter-user and CSI-beamforming dependencies over structured input and output sequences can be jointly captured. The proposed KDL-Transformer achieves substantial performance gains over both black-box learning and mathematical optimization methods.
The main contributions can be summarized as follows.
  • We propose a novel PASS-enabled downlink multi-user MISO framework, which enables the additional pinching beamforming by configuring PAs’ locations. A joint transmit and pinching beamforming optimization problem is formulated to maximize the system sum rate achieved by PASS. We propose both optimization-based and learning-based methods to tackle this problem.
  • For the optimization-based method, we propose an MM-PDD algorithm that efficiently addresses the highly nonconvex and coupled problem with strong theoretical guarantees. By reformulating the original problem into a more tractable WMMSE problem, we derive Lipschitz gradient surrogate function based on MM to handle the nonconvex complex exponential component, and then decouple the problem based on PDD. We mathematically prove that the proposed algorithm converges to stationary solutions.
@@
  • We provide numerical results to demonstrate the effectiveness of the proposed framework and algorithms, which demonstrates that: 1) The proposed PASS architecture outperforms conventional hybrid beamforming architecture even only a few PAs are activated. 2) While black-box learning algorithms are stuck at inferior points, the proposed KDL-Transformer increases over $20\%$ system performance than MM-PDD algorithm, while achieving millisecond-level fast inference on modern GPU devices.
@@
Notations: Lowercase letters (e.g., $x$) denote scalars, bold lowercase letters (e.g., $\mathbf{x}$) denote vectors, and bold uppercase letters (e.g., $\mathbf{X}$) denote matrices. $|x|$ denotes the absolute value of a real number or the modulus of a complex number. $\mathrm{Re}\left\{x\right\}$ and $\mathrm{Im}\left\{x\right\}$ denote the real and image parts of $x$, and $x^{H}$ is the complex conjugate number of $x$. $\mathbf{1}_{N\times 1}$ denotes an $N$-dimension all-ones vector. $\mathbf{X}^{\top}$ and $\mathbf{X}^{H}$ denote the transpose and the Hermitian matrix. $\left\|\mathbf{x}\right\|$ is the vector Euclidean norm, $\left\|\mathbf{X}\right\|_{\infty}=\max_{i,j}\left|x_{i,j}\right|$, and $\left\|\mathbf{X}\right\|$ is the matrix Frobenius norm.
@@ II System Model and Problem Formulation
In this paper, we propose a PASS-enabled downlink multi-user MISO framework, as shown in Fig. 1. The BS is equipped with $N$ waveguides and $M$ PAs to serve $K$ single-antenna users. Each waveguide stretches thousands of wavelengths. There are $L$ PAs flexibly pinched along each waveguide, and the total number of PAs is $M=L\times N$. By activating PAs at different points along the waveguides, both the phases of incident signals and the large-scale fading can be altered. Let $\mathcal{N}$ and $\mathcal{M}=\mathcal{L}_{1}\cup\mathcal{L}_{2}\cup\dots\mathcal{L}_{N}$ denote the sets of all the waveguides and PAs, respectively, where $\mathcal{L}_{n}=\left\{a_{n,1},a_{n,2},\dots,a_{n,L}\right\}$ is the collection of PAs at waveguide $n$. To support spatial multiplexing, each waveguide is connected to a dedicated radio frequency (RF) chain, which converts the signal multiplexed at the baseband and feeds it into the waveguide. To ensure multiplexing gains, we assume $N=K$. Both waveguides and PAs are installed at a fixed height of $h^{\mathrm{PA}}$, and the resulting PASS spans across a rectangular area with a size of $S_{\mathrm{x}}\times S_{\mathrm{y}}$ $\text{m}^{2}$. The three-dimension Cartesian coordinate of the feed point for the $n$-th waveguide is given by $\bm{\eta}_{n}^{\mathrm{W}}=\left[0,y_{n}^{\mathrm{W}},h^{\mathrm{PA}}\right]^{\top}$, where $y_{n}^{\mathrm{W}}$ is the $y$-axis coordinate of this waveguide. The location of each PA $a_{n,l}$ along the $n$-th waveguide can be defined as $\bm{\eta}_{n,l}\left(x_{n,l}\right)=\left[x_{n,l},y_{n,l},h^{\mathrm{PA}}\right]^{\top}$, where $x_{n,l}$ is the adjustable pinched location of PA $a_{n,l}$ along $x$-axis, and $y_{n,l}$ is the fixed coordinate over $y$-axis that depends on the connected waveguide, i.e., $y_{n,l}=y_{n}^{\mathrm{W}}$. Let $\mathbf{x}_{n}=\left[x_{n,1},x_{n,2},\dots,x_{n,L}\right]^{\top}\in\mathbb{R}^{L\times 1}$ stack the $x$-axis locations of PAs over the $n$-the waveguide, which satisfy $0\leqslant x_{n,1}<x_{n,2}<\dots<x_{n,L}\leqslant S_{\mathrm{x}}$, $\forall n\in\mathcal{N}$. $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\right]\in\mathbb{R}^{N\times L}$ stacks the locations of all PAs. Furthermore, the location of user $k$ is denoted by $\bm{\eta}_{k}^{\mathrm{U}}=\left[x_{k}^{\mathrm{U}},y_{k}^{\mathrm{U}},0\right]^{\top}$, $\forall k\in\mathcal{K}$.
@@ II-A Signal Radiation Model
where $\lambda_{\mathrm{W}}=\tfrac{\lambda_{f}}{n_{\mathrm{eff}}}$ is the guided wavelength with $n_{\mathrm{eff}}$ being the effective refractive index of the dielectric waveguide [14, 6]. $\kappa\triangleq\tfrac{2\pi}{\lambda_{f}}$ is the wavenumber, and $\left\|\bm{\eta}_{n,l}\left(x_{n,l}\right)-\bm{\eta}_{n}^{\mathrm{W}}\right\|=x_{n,l}$ is the distance from waveguide $n$ to PA $a_{n,l}$. Since the transmitting power is equally radiated by each PA along the waveguide, the amplitude of the transmitted signal is multiplied by $1/\sqrt{L}$.
@@
Let $\mathbf{h}_{k}^{H}\left(\mathbf{X}\right)\!=\!\big[\mathbf{h}_{k,1}^{H}\!\left(\mathbf{x}_{1}\right),\mathbf{h}_{k,2}^{H}\!\left(\mathbf{x}_{2}\right),\dots,\mathbf{h}_{k,N}^{H}\!\left(\mathbf{x}_{N}\right)\big]\in\mathbb{C}^{1\times M}$ stack the channel vectors from all the PAs to user $k$. Hence, the received signal at user $k$ through the channels manipulated by PAs can be compactly expressed by
$y_{k}\!=\!\underset{\text{desired signal}}{\underbrace{\overset{\text{pinching beamforming}}{\overbrace{\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\mathbf{G}\left(\mathbf{X}\right)}}\!\mathbf{d}_{k}\widetilde{s}_{k}}}\!+\!\underset{\text{multi-user interference}}{\underbrace{\sum_{k^{\prime}\neq k}\!\!\overset{\text{pinching beamforming}}{\overbrace{\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)}}\!\!\mathbf{d}_{k^{\prime}}\widetilde{s}_{k^{\prime}}}}\!+\!n_{k}.$ (6)
@@
$\mathrm{SINR}_{k}\!=\!\frac{\left|\mathbf{h}_{k}^{H}\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k}\right|^{2}}{\sum\limits_{k^{\prime}\neq k}\!\left|\mathbf{h}_{k}^{H}\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k^{\prime}}\right|^{2}\!\!+\!\sigma^{2}}\!\!=\!\!\frac{E_{k}}{\sum\limits_{k^{\prime}\neq k}I_{k,k^{\prime}}\!+\!\sigma^{2}}.$ (9)